# Thermodynamics of quantum processes: An operational framework for free energy and reversible athermality

### v1

@@ IV Dynamical Resource Theory of Athermality
$\mathrm{Cost}^{\sqrt{\varepsilon}}(\mathcal{N},\mathcal{T}^{\beta})$ $\leq\mathrm{Dist}^{1-\varepsilon}(\mathcal{N},\mathcal{T}^{\beta})+\frac{1}{2\beta}\ln\left(\frac{1}{1-\varepsilon}\right).$ (28)
The proof is given in Appendix A.7.
Based on the assistance of adaptive and nonadaptive strategies for the channel discrimination tasks (see [62, 65] for formal discussion), for an error $\varepsilon\in(0,1)$, we define the adaptive ($\rm ad$}) and nonadaptive ($\parallel$) asymptotic athermality distillation and formation rates of a resource channel $(\mathcal{N},\mathcal{T}^{\beta})$ as
$\mathscr{C}_{\mathrm{distill}}^{{\varepsilon,\rm ad}}[\mathcal{N}]$ $:=\limsup_{n\to\infty}\frac{1}{n}\,\mathrm{Distill}^{\varepsilon}_{\rm ad}\left(\mathcal{N}^{\otimes n},\left(\mathcal{T}^{\beta}\right)^{\otimes n}\right),$ (29)
$\mathscr{C}^{\varepsilon,{\rm ad}}_{\mathrm{cost}}[\mathcal{N}]$ $:=\liminf_{n\to\infty}\frac{1}{n}\,\mathrm{Cost}^{\varepsilon}_{\rm ad}\left(\mathcal{N}^{\otimes n},\left(\mathcal{T}^{\beta}\right)^{\otimes n}\right),$ (30)
@@
It directly follows from the definitions that $\mathscr{C}_{\mathrm{distill}}^{\varepsilon,\parallel}[\mathcal{N}]\leq\mathscr{C}_{\mathrm{distill}}^{\varepsilon,{\rm ad}}[\mathcal{N}]$ and $\mathscr{C}^{\varepsilon,\parallel}_{\mathrm{cost}}[\mathcal{N}]\leq\mathscr{C}^{\varepsilon,{\rm ad}}_{\mathrm{cost}}[\mathcal{N}]$, and $\mathscr{C}_{\mathrm{distill}}^{\varepsilon,\parallel}[\mathcal{N}]\leq\mathscr{C}^{\varepsilon,\parallel}_{\mathrm{cost}}[\mathcal{N}]\leq\mathscr{C}^{\varepsilon,{\rm ad}}_{\mathrm{cost}}[\mathcal{N}]$. We prove that the asymptotic distillation and formation rates are equal.
@@
The resource theory of athermality of quantum channels with GPSCs as free operations is asymptotically reversible, for any $\varepsilon\in(0,1)$,
$\mathscr{C}_{\mathrm{distill}}^{\varepsilon,\parallel}[\mathcal{N}]$ $=\mathscr{C}_{\mathrm{cost}}^{\varepsilon,\parallel}[\mathcal{N}]=\mathscr{C}_{\mathrm{distill}}^{\varepsilon,{\rm ad}}[\mathcal{N}]=\mathscr{C}_{\mathrm{cost}}^{\varepsilon,{\rm ad}}[\mathcal{N}]$
$=\frac{1}{2}D[\mathcal{N}\|\mathcal{T}^{\beta}]=\frac{\beta}{2}F^{\beta}[\mathcal{N}].$ (33)
@@
For any $\varepsilon\in(0,1)$, we have from [62, Theorem 1],
@@
and from [57, Theorem 4.1],
$\mathscr{C}_{\mathrm{cost}}^{\varepsilon,\parallel}[\mathcal{N}]=\mathscr{C}_{\mathrm{cost}}^{\varepsilon,{\rm ad}}[\mathcal{N}]=\frac{1}{2}D[\mathcal{N}\|\mathcal{T}^{\beta}].$ (35)
@@
Both the athermality distillation and formation rates of a quantum resource channel $(\mathcal{N},\mathcal{T}^{\beta})$ converge to the same value, which is half the relative entropy between the channel and the absolutely thermal channel $\mathcal{T}^{\beta}$. Therefore, the resource theory of athermality under GPSCs is reversible. The asymptotic athermality distillation and formation rates of a tensor-product quantum resource channel $(\mathcal{N}\otimes\mathcal{M},\mathcal{T}^{\beta})$, where the channel outputs of quantum channels $\mathcal{N}_{A^{\prime}\to A}$ and $\mathcal{M}_{B^{\prime}\to B}$ are noninteracting and $\mathcal{T}^{\beta}_{A^{\prime}B^{\prime}\to AB}=\mathcal{T}^{\beta}_{A^{\prime}\to A}\otimes\mathcal{T}^{\beta}_{B^{\prime}\to B}$, are equal to $\frac{\beta}{2}\left(F^{\beta}[\mathcal{N}]+F^{\beta}[\mathcal{M}]\right)$. This is a straightforward implication of the additivity of the free energy under tensor-product channels with noninteracting channel outputs.
@@ IV.1 Remarks on the resource theory of athermality
We note that in [19], the restricted form of the free energy associated to athermal channel, where reference to the channel are not considered, is shown to be related to the maximal work distillation. In particular, they showed that for the asymptotically many uses of each channel, the distillable work is an additive function of the considered channels. The allowed free operations for their task are thermal operations, a strict subset of Gibbs preserving channels. Our approach here can be thought of as a generalization where access to a reference to the channel is allowed along with adaptive strategies for the uses of the channel in the distillation (and formation) tasks. The access to the channel reference and adaptive strategy for distillation (and formation) tasks provide a general approach in resource-theoretic framework that involves channel discrimination. The rates of distillation and formation will never be less than that of the restricted case with parallel uses of channels and no allowance of the reference to the channel. The sandwiched Rényi free energy is additive for tensor-product channels with noninteracting outputs for $\alpha\in[\frac{1}{2},\infty)$, including the free energy ($\alpha=1$). This additivity property of the free energy makes the asymptotic rates of distillation and formation additive in our resource theory of athermality.
@@ V Informational aspects of free energy
The following corollary is a direct from the above proposition.
@@ VI Thermodynamical free energy and work extraction
For a quantum state $\rho_{A}$, the entropy $S(\rho)$ is $S(\rho):=S(A)_{\rho}:=-D(\rho_{A}\|\mathbbm{1}_{A})$ and the energy $E(\rho)$ is $E(\rho)=\langle\widehat{H}_{A}\rangle_{\rho}=\operatorname{tr}[\widehat{H}_{A}\rho_{A}]$. For a fixed inverse temperature $\beta$, the thermodynamic non-equilibrium free energy $F^{\beta}_{c}(\rho)$ of a quantum state $\rho_{A}$ is related to its entropy and energy as
$F^{\beta}_{c}(\rho)=E(\rho)-\beta^{-1}S(\rho)=\beta^{-1}D(\rho\|\widehat{\gamma}^{\beta}),$ (52)
@@
Definition 3 (Thermodynamic free energy).
The generalized thermodynamic free energy $\mathbf{F}^{\beta}_{c}[\mathcal{N}]$ of an arbitrary quantum channel $\mathcal{N}_{A^{\prime}\to A}$ is defined as
$\mathbf{F}^{\beta}_{c}[\mathcal{N}]:=\beta^{-1}\mathbf{D}[\mathcal{N}\|\widehat{\mathcal{T}}^{\beta}],$ (54)
where $\widehat{\mathcal{T}}^{\beta}=\mathcal{R}^{\widehat{\gamma}^{\beta}}$, i.e., $\widehat{\mathcal{T}}^{\beta}(\rho_{A^{\prime}})=\operatorname{tr}[\rho_{A^{\prime}}]\widehat{\gamma}^{\beta}_{A}$ is a replacer map that is completely positive but not trace-preserving. The thermodynamic non-equilibrium free energy $F^{\beta}_{c}[\mathcal{N}]$ of a quantum channel $\mathcal{N}_{A^{\prime}\to A}$ is $F^{\beta}_{c}[\mathcal{N}]=\beta^{-1}D[\mathcal{N}\|\widehat{\mathcal{T}}^{\beta}]$, obtained by taking generalized channel divergence to be relative entropy.
The function $\mathbf{D}[\mathcal{N}\|\widehat{\mathcal{T}}^{\beta}]=-\mathbf{S}^{\beta}[\mathcal{N}]$, where $\mathbf{S}^{\beta}[\mathcal{N}]$ is called the generalized thermal entropy of the channel $\mathcal{N}$ [25]. We see that $\mathcal{F}^{\beta}_{c}[\mathcal{N}]=\beta^{-1}\mathcal{S}^{\beta}[\mathcal{N}]$. The thermodynamic sandwiched Rényi free divergence $F^{\beta}_{c,\alpha}[\mathcal{N}]=\beta^{-1}D_{\alpha}[\mathcal{N}\|\widehat{T}^{\beta}]$ for $\alpha\in[\frac{1}{2},1)\cup(1,\infty]$. The sandwiched Rényi free energy of the absolutely thermal channel reduces to the Helmholtz free energy,
$F^{\beta}_{c,\alpha}[\widehat{T}^{\beta}]=-\ln Z^{\beta}_{A},$ (55)
@@
$F_{c,\alpha}[\mathcal{N}]=-\beta^{-1}S_{\alpha}[\mathcal{N}].$ (56)
@@
$F^{\beta}_{\alpha}[\mathcal{N}]=F^{\beta}_{c,\alpha}[\mathcal{N}]+\beta^{-1}\ln Z^{\beta}_{A},$ (57)
for $\alpha\in[\frac{1}{2},\infty]$. $F^{\beta}_{c}[\mathcal{N}]$ satisfies all the axiomatic properties (A1-A6) as desired from the free energy function. We also have
@@
where it suffices to optimize only over pure states $\psi_{RA^{\prime}}$. The thermodynamic sandwiched Rényi free divergence $F_{c,\alpha}[\mathcal{N}]$, for $\alpha\in[\frac{1}{2},1)\cup(1,\infty)$ satisfies the axiomatic properties (A1A5)(A1-A5) and quasi-convexity.
The following lemma follows from the definitions of the resource-theoretic and thermodynamic free energies.
@@
$F^{\beta}_{c,\alpha}[\mathcal{N}]-F^{\beta}_{c,\alpha}[\mathcal{T}^{\beta}]=F^{\beta}_{\alpha}[\mathcal{N}].$ (59)
@@
The thermodynamic free energy $F^{\beta}_{c}[\mathcal{N}]$ of a quantum channel $\mathcal{N}_{A^{\prime}\to A}$ is upper bounded as, when the reference $R$ is noninteracting with AA ($\widehat{H}_{RA}=\widehat{H}_{R}\otimes\mathbbm{1}_{A}+\mathbbm{1}_{R}\otimes\widehat{H}_{A}$),
$F^{\beta}_{c}[\mathcal{N}]\leq E[\mathcal{N}]-\beta^{-1}S[\mathcal{N}].$ (64)
The bound is saturated for the replacer channels $\mathcal{R}^{\omega}_{A^{\prime}\to A}$, for $\omega\in\operatorname{St}(A)$ and $\widehat{H}^{\rm int}_{RA}=0$, $F^{\beta}_{c}[\mathcal{R}^{\omega}]=E[\mathcal{R}^{\omega}]-\beta^{-1}S[\mathcal{R}^{\omega}]$.
@@
and $F^{\beta}[\mathcal{N}]=F^{\beta}_{c}[\mathcal{N}]+\beta^{-1}\ln Z^{\beta}_{A}$.
@@
$F^{\beta}_{c}[\mathcal{N}]$ $=\sup_{\psi\in\operatorname{St}(RA^{\prime})}\left[E(\mathcal{N}(\psi_{RA^{\prime}}))-E(\psi_{R})\right$.
@@
We conclude the proof by observing that for a replacer channel $\mathcal{R}^{\omega}$, $F^{\beta}_{c}[\mathcal{R}^{\omega}]=F^{\beta}_{c}(\omega)=E(\omega)-\beta^{-1}S(\omega)=E[\mathcal{R}^{\omega}]-\beta^{-1}S[\mathcal{R}^{\omega}]$, using the fact that $R$ and $A$ are noninteracting. ∎
@@
The generalized resource-theoretic free energy $\mathbf{F}^{\beta}[\mathcal{N}]$ and the generalized thermodynamic free energy $\mathbf{F}^{\beta}_{c}[\mathcal{N}]$ of an arbitrary quantum channel $\mathcal{N}_{A^{\prime}\to A}$ upper bounds the generalized resource-theoretic free energy $\mathbf{F}^{\beta}(\Phi^{\mathcal{N}}_{RA})$ and the generalized thermodynamic free energy $\mathbf{F}^{\beta}_{c}(\Phi^{\mathcal{N}}_{RA})$ of its Choi state $\Phi^{\mathcal{N}}_{RA}$, respectively,
@@
$\mathbf{F}^{\beta}_{c}[\mathcal{N}]$ $\geq\mathbf{F}^{\beta}_{c}(\Phi^{\mathcal{N}}_{RA})+\beta^{-1}\ln|A^{\prime}|,$ (67)

### v2

@@ IV Dynamical Resource Theory of Athermality
$\mathrm{Cost}^{\sqrt{\varepsilon}}(\mathcal{N},\mathcal{T}^{\beta})$ $\leq\mathrm{Dist}^{1-\varepsilon}(\mathcal{N},\mathcal{T}^{\beta})+\frac{1}{2}\ln\left(\frac{1}{1-\varepsilon}\right).$ (28)
Based on the assistance of adaptive and nonadaptive strategies for the channel discrimination tasks (see [65, 28, 68] for formal discussion), for an error $\varepsilon\in(0,1)$, we define the adaptive ($\rm ad$}) and nonadaptive ($\parallel$) asymptotic athermality distillation and formation rates of a resource channel $(\mathcal{N},\mathcal{T}^{\beta})$ as
$\mathscr{C}_{\mathrm{distill}}^{{\varepsilon,\rm ad}}[\mathcal{N}]$ $:=\limsup_{n\to\infty}\frac{1}{n}\,\mathrm{Distill}^{\varepsilon}_{\rm ad}\left(\mathcal{N}^{(n)},\left(\mathcal{T}^{\beta}\right)^{(n)}\right),$ (29)
$\mathscr{C}^{\varepsilon,{\rm ad}}_{\mathrm{cost}}[\mathcal{N}]$ $:=\liminf_{n\to\infty}\frac{1}{n}\,\mathrm{Cost}^{\varepsilon}_{\rm ad}\left(\mathcal{N}^{(n)},\left(\mathcal{T}^{\beta}\right)^{(n)}\right),$ (30)
@@
It directly follows from the definitions that $\mathscr{C}_{\mathrm{distill}}^{\varepsilon,\parallel}[\mathcal{N}]\leq\mathscr{C}_{\mathrm{distill}}^{\varepsilon,{\rm ad}}[\mathcal{N}]$ and $\mathscr{C}^{\varepsilon,\parallel}_{\mathrm{cost}}[\mathcal{N}]\leq\mathscr{C}^{\varepsilon,{\rm ad}}_{\mathrm{cost}}[\mathcal{N}]$, and $\mathscr{C}_{\mathrm{distill}}^{\varepsilon,\parallel}[\mathcal{N}]\leq\mathscr{C}^{\varepsilon,\parallel}_{\mathrm{cost}}[\mathcal{N}]\leq\mathscr{C}^{\varepsilon,{\rm ad}}_{\mathrm{cost}}[\mathcal{N}]$. We prove that the asymptotic distillation and formation rates are equal under nonadaptive strategies.
@@
The resource theory of athermality of quantum channels with GPSCs as free operations is asymptotically reversible under parallel uses of channels, i.e., for any $\varepsilon\in(0,1)$ and an arbitrary (square) quantum resource channel $(\mathcal{N},\mathcal{T}^{\beta})$,
$\mathscr{C}_{\mathrm{distill}}^{\varepsilon,\parallel}[\mathcal{N}]=\mathscr{C}_{\mathrm{distill}}^{\varepsilon,{\rm ad}}[\mathcal{N}]=\mathscr{C}_{\mathrm{cost}}^{\varepsilon,\parallel}[\mathcal{N}]$ $=\frac{1}{2}D[\mathcal{N}\|\mathcal{T}^{\beta}]$
$=\frac{\beta}{2}F^{\beta}[\mathcal{N}].$ (33)
@@
For any $\varepsilon\in(0,1)$, we have from [65, Theorem 1],
@@
and from [58, Theorem 4.1],
$\mathscr{C}_{\mathrm{cost}}^{\varepsilon,\parallel}[\mathcal{N}]=\frac{1}{2}D[\mathcal{N}\|\mathcal{T}^{\beta}].$ (35)
@@
It follows from [65, Theorem 2] that for any athermality distillation scheme which seeks to distill at a rate strictly higher than $\frac{1}{2}D[\mathcal{N}\|\mathcal{T}^{\beta}]$, the probability of successful distillation decays to zero exponentially fast with the number of channel uses of $\mathcal{N}$. That is, $\frac{1}{2}D[\mathcal{N}\|\mathcal{T}^{\beta}]$ also serves the strong converse bound on the asymptotic athermality distillation rate of the channel $\mathcal{N}$, irrespective of whether the strategy used is adaptive or nonadaptive.
Under parallel uses of channels, both the athermality distillation and formation rates of a quantum resource channel $(\mathcal{N},\mathcal{T}^{\beta})$ converge to the same value, which is half the relative entropy between the channel and the absolutely thermal channel $\mathcal{T}^{\beta}$. Therefore, the resource theory of athermality under GPSCs is reversible. The asymptotic athermality distillation and formation rates of a tensor-product quantum resource channel $(\mathcal{N}\otimes\mathcal{M},\mathcal{T}^{\beta})$, where the channel outputs of quantum channels $\mathcal{N}_{A^{\prime}\to A}$ and $\mathcal{M}_{B^{\prime}\to B}$ are noninteracting and $\mathcal{T}^{\beta}_{A^{\prime}B^{\prime}\to AB}=\mathcal{T}^{\beta}_{A^{\prime}\to A}\otimes\mathcal{T}^{\beta}_{B^{\prime}\to B}$, are equal to $\frac{\beta}{2}\left(F^{\beta}[\mathcal{N}]+F^{\beta}[\mathcal{M}]\right)$. This is a straightforward implication of the additivity of the free energy under tensor-product channels with noninteracting channel outputs.
@@ IV.1 Remarks on the resource theory of athermality
We note that in [19], the restricted form of the free energy associated to athermal channel, where reference to the channel are not considered, is shown to be related to the work distillation. The allowed free operations for their distillation tasks are thermal operations, a strict subset of Gibbs preserving channels. They showed that for the asymptotically many uses of each channel, the distillable work is an additive function of the considered channels. In [19, 70], the identity channel has zero distillable work which is in stark contrast with our framework where the identity channel is one of the most resourceful athermal channel. We take an alternate approach where access to a reference to the channel is allowed along with adaptive strategies for the uses of the channel in the distillation tasks. The access to the channel reference and adaptive strategy for distillation tasks provide a general approach in resource-theoretic framework that involves channel discrimination. The rates of distillation will never be less than that of the restricted case with parallel uses of channels and no allowance of the reference to the channel. The sandwiched Rényi free energy is additive for tensor-product channels with noninteracting outputs for $\alpha\in[\frac{1}{2},\infty)$, including the free energy ($\alpha=1$). This additivity property of the free energy makes the asymptotic rates of distillation additive in our resource theory of athermality.
@@ V Informational aspects of free energy
The corollary given below follows directly from the above proposition.
@@ VI Thermodynamical free energy and work extraction
For a quantum state $\rho_{A}$, the entropy $S(\rho)$ is $S(\rho):=S(A)_{\rho}:=-D(\rho_{A}\|\mathbbm{1}_{A})$ and the energy $E(\rho)$ is $E(\rho)=\langle\widehat{H}_{A}\rangle_{\rho}=\operatorname{tr}[\widehat{H}_{A}\rho_{A}]$. For a fixed inverse temperature $\beta$, the thermodynamic non-equilibrium free energy $F^{\beta}_{\rm T}(\rho)$ of a quantum state $\rho_{A}$ is related to its entropy and energy as
$F^{\beta}_{\rm T}(\rho)=E(\rho)-\beta^{-1}S(\rho)=\beta^{-1}D(\rho\|\widehat{\gamma}^{\beta}),$ (52)
@@
Definition 3 (Thermal free energy).
The generalized thermal free energy $\mathbf{F}^{\beta}_{\rm T}[\mathcal{N}]$ of an arbitrary quantum channel $\mathcal{N}_{A^{\prime}\to A}$ is defined as
$\mathbf{F}^{\beta}_{\rm T}[\mathcal{N}]:=\beta^{-1}\mathbf{D}[\mathcal{N}\|\widehat{\mathcal{T}}^{\beta}],$ (54)
where $\widehat{\mathcal{T}}^{\beta}=\mathcal{R}^{\widehat{\gamma}^{\beta}}$, i.e., $\widehat{\mathcal{T}}^{\beta}(\rho_{A^{\prime}})=\operatorname{tr}[\rho_{A^{\prime}}]\widehat{\gamma}^{\beta}_{A}$ is a replacer map that is completely positive but not trace-preserving. The thermodynamic non-equilibrium free energy $F^{\beta}_{\rm T}[\mathcal{N}]$ of a quantum channel $\mathcal{N}_{A^{\prime}\to A}$ is $F^{\beta}_{\rm T}[\mathcal{N}]=\beta^{-1}D[\mathcal{N}\|\widehat{\mathcal{T}}^{\beta}]$, obtained by taking generalized channel divergence to be relative entropy.
The function $\mathbf{D}[\mathcal{N}\|\widehat{\mathcal{T}}^{\beta}]=-\mathbf{S}^{\beta}[\mathcal{N}]$, where $\mathbf{S}^{\beta}[\mathcal{N}]$ is called the generalized thermal entropy of the channel $\mathcal{N}$ [25]. We see that $\mathcal{F}^{\beta}_{\rm T}[\mathcal{N}]=\beta^{-1}\mathcal{S}^{\beta}[\mathcal{N}]$. The thermodynamic sandwiched Rényi free divergence $F^{\beta}_{{\rm T},\alpha}[\mathcal{N}]=\beta^{-1}D_{\alpha}[\mathcal{N}\|\widehat{\mathcal{T}}^{\beta}]$ for $\alpha\in[\frac{1}{2},1)\cup(1,\infty]$. The sandwiched Rényi free energy of the absolutely thermal channel reduces to the Helmholtz free energy,
$F^{\beta}_{{\rm T},\alpha}[\widehat{\mathcal{T}}^{\beta}]=-\beta^{-1}\ln Z^{\beta}_{A},$ (55)
@@
$F^{\beta}_{{\rm T},\alpha}[\mathcal{N}]=-\beta^{-1}S_{\alpha}[\mathcal{N}].$ (56)
@@
$F^{\beta}_{\alpha}[\mathcal{N}]=F^{\beta}_{{\rm T},\alpha}[\mathcal{N}]+\beta^{-1}\ln Z^{\beta}_{A},$ (57)
for $\alpha\in[\frac{1}{2},\infty]$. $F^{\beta}_{\rm T}[\mathcal{N}]$ satisfies all the axiomatic properties (A1-A6) as desired from the free energy function. We also have
@@
where it suffices to optimize only over pure states $\psi_{RA^{\prime}}$. The thermodynamic sandwiched Rényi free divergence $F_{{\rm T},\alpha}[\mathcal{N}]$, for $\alpha\in[\frac{1}{2},1)\cup(1,\infty)$ satisfies the axiomatic properties (A1A5)(A1-A5) and quasi-convexity.
The following lemma follows from the definitions of the resource-theoretic and thermal free energies.
@@
$F^{\beta}_{{\rm T},\alpha}[\mathcal{N}]-F^{\beta}_{{\rm T},\alpha}[\mathcal{T}^{\beta}]=F^{\beta}_{\alpha}[\mathcal{N}].$ (59)
@@
The thermal free energy $F^{\beta}_{\rm T}[\mathcal{N}]$ of a quantum channel $\mathcal{N}_{A^{\prime}\to A}$ is upper bounded as, when the reference $R$ is noninteracting with AA ($\widehat{H}_{RA}=\widehat{H}_{R}\otimes\mathbbm{1}_{A}+\mathbbm{1}_{R}\otimes\widehat{H}_{A}$),
$F^{\beta}_{\rm T}[\mathcal{N}]\leq E[\mathcal{N}]-\beta^{-1}S[\mathcal{N}].$ (64)
The bound is saturated for the replacer channels $\mathcal{R}^{\omega}_{A^{\prime}\to A}$, for $\omega\in\operatorname{St}(A)$ and $\widehat{H}^{\rm int}_{RA}=0$, $F^{\beta}_{\rm T}[\mathcal{R}^{\omega}]=E[\mathcal{R}^{\omega}]-\beta^{-1}S[\mathcal{R}^{\omega}]$.
@@
and $F^{\beta}[\mathcal{N}]=F^{\beta}_{\rm T}[\mathcal{N}]+\beta^{-1}\ln Z^{\beta}_{A}$.
@@
$F^{\beta}_{\rm T}[\mathcal{N}]$ $=\sup_{\psi\in\operatorname{St}(RA^{\prime})}\left[E(\mathcal{N}(\psi_{RA^{\prime}}))-E(\psi_{R})\right$.
@@
We conclude the proof by observing that for a replacer channel $\mathcal{R}^{\omega}$, $F^{\beta}_{\rm T}[\mathcal{R}^{\omega}]=F^{\beta}_{\rm T}(\omega)=E(\omega)-\beta^{-1}S(\omega)=E[\mathcal{R}^{\omega}]-\beta^{-1}S[\mathcal{R}^{\omega}]$, using the fact that $R$ and $A$ are noninteracting. ∎
@@
The generalized resource-theoretic free energy $\mathbf{F}^{\beta}[\mathcal{N}]$ and the generalized thermal free energy $\mathbf{F}^{\beta}_{\rm T}[\mathcal{N}]$ of an arbitrary quantum channel $\mathcal{N}_{A^{\prime}\to A}$ upper bounds the generalized resource-theoretic free energy $\mathbf{F}^{\beta}(\Phi^{\mathcal{N}}_{RA})$ and the generalized thermal free energy $\mathbf{F}^{\beta}_{\rm T}(\Phi^{\mathcal{N}}_{RA})$ of its Choi state $\Phi^{\mathcal{N}}_{RA}$, respectively,
@@
$\mathbf{F}^{\beta}_{\rm T}[\mathcal{N}]$ $\geq\mathbf{F}^{\beta}_{\rm T}(\Phi^{\mathcal{N}}_{RA})+\beta^{-1}\ln|A^{\prime}|,$ (67)